Midnight Sunset at "Camp 2" -- 12 May 2013

Thursday, May 2, 2013

work progressing, weather closing

Yesterday afternoon (1 May), I got a call from Dirk van As giving status update for the team...

weather stations LAN_U and KAN_M are/were running fine

3rd ~20 m core being drilled today 2 May

Liam and Mark returning to DYE-2 to return snow drift sensors and resupply with fuel.

After much effort, got ground penetrating radar (GPR) working

Dirk had to dig 1.7 m down to service a GPS for A. Hubbard.

weather should be closing in on them today 2 May.

The team plans to return to DYE-2 for a crew rotation 6 or 7 May

Dirk aims to return to Copenhagen 10 May

No comments:

Post a Comment